OpenStudy (anonymous):

A car radiator contains 8 liters of a 32% antifreeze solution. How many liters of this solution should be drained and replaced with pure antifreeze to get a 49% solution?

OpenStudy (anonymous):

OKay so for this... damn it i have to go okay sorry bye!!

OpenStudy (anonymous):

8 liters

OpenStudy (anonymous):

is 4L 2L 3L or 1.75

OpenStudy (anonymous):

drain 8 liters of 32% and replace 8 liters of 49%

OpenStudy (dumbcow):

haha thats the lazy way :)

OpenStudy (anonymous):

my only choices are 2 3 4 or 1.75 not 8

OpenStudy (dumbcow):

system of 2 equations Let x be 32% antifreeze, y is amount drained and replaced x+y = 8 .32x + y = .49

OpenStudy (dumbcow):

oops it should be 8*.49 use elimination subtract row 1 - row 2 x + y = 8 -.32x -y = -(8*.49) __________________ .64x +0 = 4.08 x = 4.08/.64 x = 6.375 y = 8-x y = 8-6.375 y=1.625 1.625L should be drained and replaced

OpenStudy (dumbcow):

wierd i did something really wrong sorry answer is 2L